Solve by elimination: 5x -6y = -32
3x+6y= 48



Answer by oscargut(686)

5x -6y = -32
3x+6y = 48
8x = 16
x=2
using first eq
5(2)-6y=-32
-6y=-42
y=7
Answer: x=2,y=7